# Double-quantum-dot Andreev molecules: Phase diagrams and critical evaluation of effective models

### v1

@@ III.2 Failure of the ZBW approximation
However, going further to the case of $\Gamma/U=0.05$, reveals ZBW to fail completely even in qualitative terms, as triplet GSs are not predicted by ZBW approximation at all for a broad range of $\Delta/U$ spanning approximately the interval of $<10^{-3},10^{-1}>$. On the other hand, the shape of the singlet and doublet regions away from the central portion of the phase diagram seem to match the exact results if appropriate values of $\Gamma^{\mathrm{ZBW}}_{0}$ are selected. For instance, at $\Delta/U=5\times 10^{-3}$, there seem to be a relatively good match between Figs. 7(c3) and (c7) in terms of singlet and doublet phases.
@@ III.3 Extended ZBW model
Significant lead-mediated interactions present in the parallel DQD setup generate triplet GSs across a broad parameter range, challenging both AL-based models and the ZBW approximation. To address this, we extended the ZBW model in Sec. I.4 by including an additional SC orbital. The resulting eZBW model introduces then a new effective parameter, the hybridization term $V^{\mathrm{ZBW}}_{1}$, which couples orbital $1$ to the ZBW orbital labeled 00.
Consequently, applying eZBW involves handling a multi-parametric model with unknown effective values. In order to manage such a challenge, we first assume $V^{\mathrm{ZBW}}_{1}$ to be the same for all level detunings $\delta_{1}$ and $\delta_{2}$, while $\Delta$, $U_{1}$ and $U_{2}$ remain equal to their bare values as before. This leaves $V^{\mathrm{ZBW}}_{0}$, $V^{\mathrm{ZBW}}_{1}$, and possibly $t$, to be adjusted within eZBW. Unlike in the case of the AL and ZBW, we show that the additional degree of freedom brought in by the extra orbital allows to qualitatively model even the presence of triplet GSs. However, we emphasize that the presented analysis should be understood as a preliminary one. Establishing eZBW as a universal effective model for a present scenario requires additional research focused on ABS spectroscopy and other static properties, areas that extend, however, far beyond the scope of this paper.
We stress however that our analysis should be understood as a preliminary one, further research should therefore concentrate on the ABS spectroscopy and other static properties. However, such an analysis is largely beyond the scope of the present paper.
@@
Despite many simplifying assumptions about the parameters, the comprehensive analysis of the eZBW model in regards of only the GSs remains still tedious. However, to assess the potential improvements over other effective models, it is particularly important to focus especially on regimes where both the (G)AL and ZBW models have previously failed. Such regimes are observed when lead-mediated interactions dominate the system, which happens especially as $\nu\neq 1$. We therefore select the coupling-symmetric scenario at $\nu=1$, while identical dots ($U_{1}=U_{2}=U$) are considered without inter-dot interaction. To demonstrate the possibilities of eZBW we further specifically address the DQD parameters of Fig. 2 with $\Gamma/U=0.05$ and select two cases: $\Delta/U=5\times 10^{-4}$ and $\Delta/U=5\times 10^{-2}$, which have previously been shown as inaccessible by ZBW. To fully explore the eZBW parametric space, we then vary $V^{\mathrm{ZBW}}_{0}/U$ and $V^{\mathrm{ZBW}}_{1}/U$ across several orders of magnitude. Since the resulting data is rather elaborate, we show the complete results in the Appendix in Figs. 10 and 11 with panels ordered by increasing $V^{\mathrm{ZBW}}_{0}$ horizontally and $V^{\mathrm{ZBW}}_{1}$ vertically.
@@
Starting with the $\Delta/U=5\times 10^{-2}$ case, we notice that a significant portion of the eZBW-parametric space matches the the exact phase diagram observed in Fig. 2(f) quite well. Selecting Fig. 10(c3) as the best fit, yields then $V^{\mathrm{ZBW}}_{0}/U=0.0125$ and $V^{\mathrm{ZBW}}_{1}/U=0.02$. Both hybridization terms are thus of similar magnitude, with $V^{\mathrm{ZBW}}_{0}$ giving a tunneling rate of the order of $\Gamma^{\mathrm{ZBW}}_{0}/U\approx 10^{-4}$. Reducing then the BCS gap further to $\Delta/U=5\times 10^{-4}$, once again positively demonstrates eZBW’s capabilities in reproducing the exact phase diagram shown in Fig. 2(d). However, the effective tunneling rates inserted into the eZBW drop even further to $\Gamma^{\mathrm{ZBW}}_{0}/U\approx 10^{-6}$ as Fig. 11(d3) with $V^{\mathrm{ZBW}}_{0}/U=0.001$ and $V^{\mathrm{ZBW}}_{1}/U=5\times 10^{-5}$ represents the best fit to the exact result of Fig. 3(g).
The analysis of the two cases shows that determining both $V^{\mathrm{ZBW}}_{0}$ and $V^{\mathrm{ZBW}}_{1}$ is relatively costly even at $t=0$. To avoid such problems we now try to set both equal to each other which was approximatively satisfied in the first case. Moreover, we fix their values now to $V^{\mathrm{ZBW}}_{0}/U=V^{\mathrm{ZBW}}_{1}/U=0.1$ and change the ratio $\Delta/U$ in analogy to Fig. 2 with resulting phase diagrams shown in Fig. 9. Since the effective values of $V^{\mathrm{ZBW}}_{0}/U$ and $V^{\mathrm{ZBW}}_{1}/U$ undergo corrections of several magnitudes as $\Delta/U$ is decreased, we expect this simple approach to break down as $\Delta/U\rightarrow 1$. Indeed, for $\Delta=0$, a regime dominated by Kondo effects, the simplified approach to eZBW incorrectly predicts a triplet GS. However, as the gap is further increased, the triplet GS evolution becomes qualitatively comparable to the exact results, albeit larger values of $\Delta/U$ have to be matched. The erroneous predictions of eZBW in Figs. 9(a) and (b) are, of course, the artifact of selecting fixed values of hybridization terms for all values of the CS gap parameter.
Overall, eZBW offers significant improvements over other approximations. However, the rapid evolution of effective tunneling rates as $\Delta/U$ decreases necessitates extensive parametric scans, diminishing eZBW’s practical benefits. Approximate prescriptions for effective hybridization terms $V^{\mathrm{ZBW}}_{0}$ and $V^{\mathrm{ZBW}}_{1}$ have the potential to largely mitigate this disadvantage in a wide range of DQD’s parameters. The initial analysis performed based on GS properties, however, demonstrates that eZBW has the potential to outperform AL-based and ZBW effective models, which both require a large gap and/or large inter-dot hopping in order to model the present case correctly. Further research involving also ABS spectroscopy and thermodynamic quantities should thus be undertaken.
@@ V.1 Particle-hole transformations
[…]

### v2

@@ III.2 Failure of the ZBW approximation
However, going further to the case of $\Gamma/U=0.05$, reveals ZBW to fail completely even in qualitative terms, as triplet GSs are not predicted by ZBW approximation at all for a broad range of $\Delta/U$ spanning approximately the interval of $\left[10^{-3},10^{-1}\right]$. On the other hand, the shape of the singlet and doublet regions away from the central portion of the phase diagram seem to match the exact results if appropriate values of $\Gamma^{\mathrm{ZBW}}_{0}$ are selected. For instance, at $\Delta/U=5\times 10^{-3}$, there seem to be a relatively good match between Figs. 7(c3) and (c7) in terms of singlet and doublet phases.
@@ III.3 Extended ZBW model
In previous sections we have demonstrated that both AL-based models and the ZBW approximation can lead to unphysical predictions for the parallel DQD systems. We therefore emphasize that a good practice is to always benchmark simplified descriptions of QD-based problems with exact approaches like NRG when extending them beyond their proven applicability domain. On the other hand, simplified models offer valuable tools when their limitations are respected. For example, even when constrained to a specific regime, these models can provide analytical insights that aid in the design of new devices, as illustrated by the case of equilibrium quartet superconductivity [44].
As an example of such a simplified approach, we now aim to extend the ZBW model to capture the singlet-triplet transitions identified in Sec, V.2 by exact NRG calculations. Our focus lies in this particular aspect, and we stress that no rigorous claims are made regarding the broader applicability beyond the current model or the effect we aim to describe. Given the importance of lead-mediated interactions in the DQD system, it is natural to extend the ZBW approximation defined in Sec. I.4 by adding another SC orbital labeled $1$. It is supposed to couple via a new effective parameter, the hybridization term $V^{\mathrm{ZBW}}_{1}$, only to the original ZBW orbital labeled 0. We refer to this approximation as the extended ZBW (eZBW) model.
In eZBW, the parametric space has been further enlarged. In order to manage such a challenge, we first assume $V^{\mathrm{ZBW}}_{1}$ to be the same for all level detunings $\delta_{1}$ and $\delta_{2}$, while $\Delta$, $U_{1}$ and $U_{2}$ remain equal to their bare values as before. This leaves $V^{\mathrm{ZBW}}_{0}$, $V^{\mathrm{ZBW}}_{1}$, and possibly $t$, to be adjusted within eZBW. Despite these many simplifying assumptions, analysis based on eZBW remains still multi-parametric [^3] The initial values might be inputed from the corresponding surrogate calculations [86, 87]. However, further adjustments are still required..
@@
To describe the challenging scenario of singlet-triplet transitions, we select the coupling-symmetric scenario at $\nu=1$ and identical dots ($U_{1}=U_{2}=U$) without inter-dot interaction. We then address the DQD parameters of Fig. 2 with $\Gamma/U=0.05$ and select two cases: $\Delta/U=5\times 10^{-4}$ and $\Delta/U=5\times 10^{-2}$, which have previously been inaccessible by ZBW or AL-based models. We then vary $V^{\mathrm{ZBW}}_{0}/U$ and $V^{\mathrm{ZBW}}_{1}/U$ across several orders of magnitude to analyze the corresponding phase diagrams. Since the resulting data is rather elaborate, we show the complete results in the Appendix in Figs. 10 and 11 with panels ordered by increasing $V^{\mathrm{ZBW}}_{0}$ horizontally and $V^{\mathrm{ZBW}}_{1}$ vertically.
@@
Starting with the $\Delta/U=5\times 10^{-2}$ case, we notice that a significant portion of the eZBW-parametric space matches the exact phase diagram observed in Fig. 2(f) quite well. Selecting Fig. 10(c3) as the best fit yields then $V^{\mathrm{ZBW}}_{0}/U=0.0125$ and $V^{\mathrm{ZBW}}_{1}/U=0.02$. Both hybridization terms are thus of similar magnitude, with $V^{\mathrm{ZBW}}_{0}$ giving a tunneling rate of the order of $\Gamma^{\mathrm{ZBW}}_{0}/U\approx 10^{-4}$. Reducing then the BCS gap further to $\Delta/U=5\times 10^{-4}$ once again positively demonstrates eZBW’s capabilities in reproducing the exact phase diagram shown in Fig. 2(d). However, the effective tunneling rates inserted into the eZBW drop even further to $\Gamma^{\mathrm{ZBW}}_{0}/U\approx 10^{-6}$ as Fig. 11(d3) with $V^{\mathrm{ZBW}}_{0}/U=0.001$ and $V^{\mathrm{ZBW}}_{1}/U=5\times 10^{-5}$ represents the best fit to the exact result of Fig. 3(g).
The analysis of the two cases shows that determining both $V^{\mathrm{ZBW}}_{0}$ and $V^{\mathrm{ZBW}}_{1}$ is relatively costly even at $t=0$. To avoid such problems we resort to the next simplification and assume both parameters to be equal. We fix their values to $V^{\mathrm{ZBW}}_{0}/U=V^{\mathrm{ZBW}}_{1}/U=0.1$ leaving in analogy to Fig. 2 only the ratio $\Delta/U$ to be tuned. The resulting phase diagrams are shown in Fig. 9. Due to the assumption $V^{\mathrm{ZBW}}_{0}=V^{\mathrm{ZBW}}_{1}$, unphysical behavior appears in Fig. 9(a) at $\Delta=0$ and in Fig. 9(b) at $\Delta=0.005$ where Kondo effects dominate the system. Here, the simplified eZBW incorrectly predicts a sizable triplet GS in the middle of the phase diagram. Only as $\Delta/U$ is further increased, the triplet GS evolution becomes qualitatively comparable to the exact NRG results obtained earlier, albeit larger values of $\Delta/U$ are required.
Overall, eZBW allowed us to model singlet-triplet transitions of the full problem. However, the rapid evolution of effective tunneling rates as $\Delta/U$ decreases necessitates extensive parametric scans, diminishing eZBW’s practical benefits. Approximate prescription fixing $V^{\mathrm{ZBW}}_{0}=V^{\mathrm{ZBW}}_{1}$ rectified this problem, but is qualitatively applicable only in a specific range of values illustrated in Fig. 9(c)-(f). While eZBW shows potential for simplified modeling of the present system, we emphasize that a further quantitative analysis would be required.
@@ V.1 Particle-hole transformations
$-t\sum_{\sigma}\left(d_{1\sigma}^{\dagger}d_{2\sigma}^{\vphantom{\dagger}}+\textit{h.c.}\right)-\sum_{j\mathbf{k}}V_{j,\mathbf{k}}^{\vphantom{\dagger}}\left(c^{\dagger}_{\mathbf{k}\sigma}d_{j\sigma}^{\vphantom{\dagger}}+\textit{h.c.}\right)$
$-\sum_{\mathbf{k}\sigma}\varepsilon_{\mathbf{k}}c^{\dagger}_{\mathbf{k}\sigma}c_{\mathbf{k}\sigma}^{\vphantom{\dagger}}+2\sum_{\mathbf{k}}\varepsilon_{\mathbf{k}}+\sum_{\mathbf{k}}\Delta\left(c^{\dagger}_{\mathbf{k}\uparrow}c^{\dagger}_{-\mathbf{k}\downarrow}+\textit{h.c.}\right),$